OpenStudy (anonymous):

My book wants me to factor polynomials. They want me to factor out the biggest single term i can. Would my answer be right on this question? 5xexponet2+10x=5x(x+2x)

OpenStudy (mathstudent55):

5x^2 + 10x The biggest term is 5x, but when you factor it out, this is what you get: 5x(x + 2) There is no x in the 2 term

OpenStudy (stamp):

5x^2 = 5 * x * x 10x = 2 * 5 * x They each share a *5 and a *x GCF = 5x

OpenStudy (anonymous):

Thank you so much you all helped.